OpenStudy (samsan9):

what is the radius of the circle shown below? x^2+y^2-8x-10y-8=0

OpenStudy (campbell_st):

complete the square in both x and y

OpenStudy (anonymous):

\[\left\{ x^2-8x+\left( \frac{ -8 }{ 2 } \right)^2 \right\}+\left\{ y^2-10 y+\left( \frac{ -10}{ 2 } \right) ^2 \right\}=8+\left( \frac{ -8 }{ 2 } \right)^2+\left( \frac{ -10 }{ 2 } \right)^2\] ?

OpenStudy (samsan9):

what do i do with the -8?

OpenStudy (samsan9):

sorry i just dont which one the -8 belongs too or if i have to add it to the other side

OpenStudy (anonymous):

\[\left( x-h \right)^2+\left( y-k \right)^2=r^2\]

OpenStudy (campbell_st):

add it to both sides of the equation so you have terms in x and y on the left... and a value on the right

OpenStudy (campbell_st):

so to start you have x^2 - 8x + y^2 - 10y = 8

OpenStudy (samsan9):

(x^2-8x)(y^-10y)=8

OpenStudy (campbell_st):

thats the start now complete the squares... and remember to add the same values to both sides of the equation.

OpenStudy (samsan9):

i got (4,5) as the center

OpenStudy (campbell_st):

thats correct..

OpenStudy (campbell_st):

so what is the radius..?

OpenStudy (samsan9):

7

OpenStudy (campbell_st):

thats correct * + 25 + 16 = 49